OpenStudy (anonymous):

Solve 2x – 5 = 3(2x + 5). You must show all work.

OpenStudy (anonymous):

Multiply out the right side, 3 times 2x and 5 times 3 What do you get?

OpenStudy (anonymous):

2x-5=6x+15?

OpenStudy (anonymous):

sooo then add five on the left and on the right side and i get .. 2x=6x+20?

OpenStudy (anonymous):

Idk.. ha

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Sorry.. yeh thats right.. now get the 6 to the left side

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Subtract 6x from both sides

OpenStudy (anonymous):

-4x=x+20

OpenStudy (anonymous):

No: 2x = 6x + 20 -6x -6x -------------- -4x = 20 Now divide by -4

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

I got x=-5 Thank you for being so patient with me!! and not jkust giving me the answer